# COSINE-100U: Upgrading the COSINE-100 Experiment for Enhanced Sensitivity to Low-Mass Dark Matter Detection

### v1

@@ I Introduction
Numerous astronomical observations suggest that the majority of matter in the universe consists of invisible dark matter, though its nature and interactions remain elusive Clowe et al. (2006); Aghanim et al. (2020); Bertone and Hooper (2018). Despite extensive efforts to directly detect dark matter, no definitive signals have been observed Marrodán Undagoitia and Rauch (2016); Schumann (2019). The only exception is the DAMA/LIBRA experiment, which has reported annual modulation signals in a 250 kg array of NaI(Tl) detectors Bernabei et al. (1998, 2013, 2018, 2021), indicating potential dark matter-nuclei interactions Savage et al. (2009); Ko et al. (2019). However, these results remain controversial, as no other experiments have observed similar signals Schumann (2019); Workman et al. (2022), necessitating independent verification using the same NaI(Tl) crystal target materials.
@@
The COSINE-100 experiment concluded in March 2023 after 6.4 years of stable operation. It was relocated to the newly constructed Yemilab Park et al. (2024); Kim and Lee (2024) in Korea for the next phase, COSINE-100U. During the transition, the crystal encapsulation was upgraded to improve light collection by directly attaching the crystals to photomultiplier tubes (PMTs) Choi et al. (2020). A similar technique, applied in the NEON experiment for measuring coherent neutrino-nucleus scattering at nuclear reactors Choi et al. (2023), achieved approximately a 50% improvement in light yield and demonstrated long-term stability over two years Choi et al. (2024a). These improvements in light yield are expected to lower the energy threshold, thereby improving the sensitivity of COSINE-100U to dark matter, particularly in the low-mass region. For the COSINE-100U experiment at Yemilab, all eight NaI(Tl) crystals from COSINE-100 were upgraded. In this article, we detail the detector upgrade and the resulting improved sensitivities for low-mass dark matter detection.
@@ II NaI(Tl) Crystal Encapsulation
Commercial companies typically package NaI(Tl) crystals using aluminum or copper bodies with quartz windows. The crystal surfaces are often wrapped in reflective materials, such as Teflon sheets or aluminum oxide powder, to enhance light reflection.
@@
All eight COSINE-100 crystals were grown by Alpha Spectra Inc. in collaboration with low-background research efforts of the KIMS, DM-Ice, and ANAIS collaborations Kim et al. (2015); Amaré et al. (2015); Adhikari et al. (2016). These cylindrical crystals were hermetically encased in oxygen-free copper (OFC) tubes (1.5 mm thick) with quartz windows at both ends, as exemplified by crystal-6 (C6) in Fig. 1. The lateral surfaces of each crystal were wrapped in Teflon sheets before being inserted into the OFC tubes. A 12 mm-thick quartz window was coupled to the crystal using a 1.5 mm-thick optical pad, and PMTs were attached to the quartz window using a small amount of high-viscosity optical grease.
@@
Figure 1: The encapsulation design of the COSINE-100 C6 crystal. A 4.8-inch diameter crystal is encased in OFC tubes, with an optical pad and quartz windows. The 12 mm-thick quartz window guides light from the 4.8-inch diameter crystal to the 3-inch PMTs, aided by a 45 angle surface. There are three optical interface layers between the crystal and the PMTs: a 1.5 mm optical pad, a 12 mm quartz windows, and a small amount of optical grease. The PMTs are encapsulated in OFC cases to protect them from the surrounding liquid scintillator.
@@
From the perspective of light collection efficiency, the original COSINE-100 encapsulation design had several drawbacks. The three-layer optical interface–comprising the optical pad, quartz window, and optical grease–added additional reflections, reducing light collection efficiency. Although the 12 mm-thick quartz window, with a 45 angle, guided photons from the 4.8-inch diameter crystal, it was insufficient for efficiently directing photons to the 3-inch PMTs. As shown in Fig. 1, uncovered areas resulted in photon reflection, further reducing light collection efficiency.
@@
To address these issues, we developed a novel crystal encapsulation technique that involved directly attaching the PMT to the crystal using only a 2 mm-thick optical pad Choi et al. (2020). In this method, the crystal’s diameter was matched to the 3-inch diameter of the PMTs, eliminated photon absorption in the quartz and minimized photon loss due to reflections from multiple interfaces. This modification resulted in up to 22 photoelectrons (NPE)/keV, approximately 50% higher than the light yield with the COSINE-100 crystals Adhikari et al. (2018a).
@@
In this design, the crystal and PMTs were assembled as a single unit, with the PMTs becoming integral components of the airtight crystal encapsulation. A 1 mm-thick PTFE o-ring was used to seal the neck of the PMT glass between the copper cylinder’s endcap and the back of the PMT glass, preventing any air leakage. This technique was initially applied in the NEON experiment for observing coherent neutrino-nucleus scattering during an engineering run at a nuclear reactor Choi et al. (2023). However, the design revealed some weaknesses, as liquid scintillator leakage into the detector caused a gradual decrease in crystal light yield and increase of PMT-induced noise.
@@
Since then, We have improved the crystal encapsulation by separating it into two components: an inner structure to maintain a stable coupling between the PMTs and crystal, and an outer OFC case to prevent the infiltration of external air and liquid scintillator. These improvements were successfully applied to the NEON experiment, which has been collecting stable physics data for over two years Choi et al. (2024a).
II.1 Design of COSINE-100U Encapsulation
To maximize light collection efficiency, we used 3-inch diameter crystals to match the PMT diameters in the NEON experiment Choi et al. (2023). However, COSINE-100 crystals have larger diameters than the PMTs, as summarized in table 1. To address this, we designed the crystal edges to guide light from the larger crystal diameter to the 3-inch PMTs by machining beveled edges at a 45 angle, as shown in Fig. 2 (a) for the C6 crystal. During this process, the effective area of the 3-inch PMT photocathode, typically 72 mm, was considered, and the crystal was machined to a 70 mm diameter at the ends to optimize PMT attachment.
Figure 2: An example of the inner structure for the C6 crystal, designed to maintain a stable coupling between the PMTs and crystal. (a) The design of the C6 machining for guiding light to 3-inch PMTs. (b) A PTFE structure that covers the crystal and connects it to the PMTs. (c) An illustration of the crystal-PMT assembly, including 2 mm thick optical pad for coupling.
To couple the crystal with the PMTs, we designed a 5 mm-thick PTFE structure to surround the crystal, as shown in Fig. 2 (b). This PTFE structure is horizontally separable at the center, allowing the crystal to be placed inside. Each end of PTFE structure is connected to the PMT using a PTFE ring with brass bolts, applying pressure to the optical pad between the crystal and the PMT for optical light coupling, as illustrated in Fig. 2 (c).
The outer tube, made from OFC, follows the updated design of the NEON experiment Choi et al. (2024a), but with a larger diameter and length. As shown in Fig. 3 (a), the C6 tube has a length of 660 mm, a diameter of 140 mm, and a thickness of 2 mm. The side flanges (Fig. 3 (b)) are 20 mm-thick to provide sufficient pressure for a tight seal. Two grooves were machined into each side to accommodate PTFE gaskets, as shown in Fig. 3 (c). The two OFC lids (Fig. 3 (d)), each 20 mm thick, are attached to the tube flanges using twelve brass bolts through the PTFE gaskets. Each lid contains three holes for passing one high-voltage cable and two signal cables for the anode and dynode readouts Adhikari et al. (2018c). Waterproof cable glands secure the lids, preventing the ingress of liquid scintillator and external air through the cable exit holes. The final assembly of the inner and outer structures is illustrated in Fig. 3 (e).
Figure 3: An example of the outer structure for the C6 crystal, designed to prevent the infiltration of outside air and liquid scintillator. (a) The OFC tube for the C6. (b) Side view of the OFC tube showing the flange with two grooves for tight sealing using gaskets. (c) Design of the PTFE gasket. (d) OFC lid, which couples with the flange of the tube through the PTFE gasket. (e) Illustration of the combined inner and outer structure for C6
II.2 Crystal Machining and Encapsulation
To upgrade the crystals for the COSINE-100U experiment, the existing encapsulation was first removed, and the crystal edges were machined, as shown in Fig. 4. This process was performed by a specialized machining company. A dry room with a dedicated lathe machine was prepared for removing the original encapsulation and shaping the crystal edges to improve light guidance. During machining, mineral oil was continuously poured over the crystal to prevent cracks and suppress NaI(Tl) dust in the environment. Given the highly hygroscopic nature of NaI(Tl) crystals, the machined crystals were stored in a dry storage box and quickly transferred to a low-humidity, N2 gas-flushed glovebox to avoid exposure to atmospheric moisture.
Figure 4: Machining and delivery process of the C1 crystal. (a) The COSINE-100 C1 bare crystal after the removal of the original encapsulation. (b) Machining of C1 using a lathe machine. (c) The crystal storage box to transport the machined crystal to glovebox. The storage box was filled with dehumidifiers and flushed with N2 gas to prevent moisture exposure.
The humidity level inside the glovebox was maintained at below a few tens of ppm (H2O) using a molecular sieve trap and N2 gas flushing. Before polishing the crystals, N2 gas was flushed at a high flow rate of 15 liters per minute for 2 hours, which was then reduced to 5 liters per minute to minimize radon levels inside the glovebox.
After machining, the crystals were contaminated with mineral oil, NaI(Tl) powder, and other debris, as shown in Fig. 5 (a). To clean the crystal surfaces, they were gently wiped with anhydrous ethanol. Due to the small amount of water in the ethanol, this process slightly removed the outer crystal layer. Each end of the crystal was then mirror-polished using a polishing pad and SiO2 abrasives in two-steps: first with 3 $\mu$m and then with 0.5 $\mu$m particle sizes. After polishing, all surfaces were wiped with cleanroom wipes soaked in anhydrous isopropanol, resulting in a shiny finish on the crystal surfaces, as shown in Fig. 5 (b). The stability of the polished crystal surfaces was tested in the glovebox for a week, during which no visible changes were observed.
Figure 5: Example of the COSINE-100U crystal encapsulation process. (a) The C1 crystal was delivered to a low-humidity glovebox after machining for light guidance. (b) The crystal surface was cleaned and polished. (c) The inner structure was assembled by directly attaching the PMTs to the crystal using an optical pad. (d) The inner structure was inserted into the outer OFC tubes and sealed with two OFC lids using PTFE gaskets. (e) The outer lids feature three outlets for signal and high-voltage cables, secured with waterproof cable glands.
[…]
The polished crystals were wrapped with 250 $\mu$m thick Teflon sheets to enhance light collection. They were directly attached to the PMTs using 2 mm-thick optical pads, held in place by a PTFE inner structure with six screw holes, as shown in Fig. 5 (c). PTFE pressure rings were used to apply sufficient pressure between the crystal and PMTs for proper optical coupling, secured with brass bolts, nuts, and washers. Special locking washers were used to prevent loosening due to vibrations during transport and operation. The entire assembly was then placed inside a copper case and sealed with copper lids using brass bolts and nuts, as shown in Fig. 5 (d). To prevent external contamination, PTFE gaskets and double layers of PTFE reflective sheets were inserted at the flange joints. Cables were routed through cable glands at the center of the copper lid, as shown in Fig. 5 (e).
@@ III.1 Measurement Setup
Upon assembling the COSINE-100 crystals, the Yemilab facility was not yet ready for the operation of the COSINE-100U experiment. For the initial crystal characterization, we employed a simple shielding setup at sea level in the experimental hall of the Institute for Basic Science (IBS) in Korea. This setup consisted of two layers of shielding: 10 cm-thick lead and 20 cm-thick liquid scintillator, which also functioned as an active veto detector. The liquid scintillator was housed within a 124.5 cm $\times$ 49.5 cm $\times$ 49.5 cm cubic stainless steel box. Three 8-inch PMTs were used to read the signals from the liquid scintillator. This setup was initially developed as a prototype detector for the NEOS experiment Ko et al. (2017) and was reused for this test. Inside the container, an acrylic table was used to install one NaI(Tl) crystal for the first characterization, as shown in Fig. 6.
Figure 6: (a) Schematic view of the shield structure used in the sea level measurement at the Institute for Basic Science. (b) The C1 crystal is installed in this setup for light yield and stability measurements.
@@ III.2 Light Yield Measurement
We first measured the light yield of the newly encapsulated crystal by irradiating it with 59.54 keV $\gamma$-rays from a 241Am source. The mean charge corresponding to a single photoelectron (SPE) was determined by analyzing trailing isolated cluster pulses in the 5–7 and 6–8 $\mu$s windows, which were 2.6–4.6 and 3.6–5.6 $\mu$s from the typical pulse start Adhikari et al. (2018c), to avoid multiple photoelectrons. We simultaneously fit the 5–7 and 6–8 $\mu$s clusters with models of up to four photoelectrons clusters Choi et al. (2024b, a), as shown in Fig. 7. The number of photoelectrons (NPE) was calculated by dividing the integrated charge of the main pulse (within 5 $\mu$s of the pulse start) by the charge of SPE.
@@
Figure 7: The charge distribution of the isolated clusters in the 5–7 $\mu$s (red dots) and 6–8 $\mu$s (blue dots) windows. The distribution is modeled with up to four NPE clusters, simultaneously fitted for both the 5–7 and 6–8 $\mu$s windows, yielding a mean charge of 597.2 ADC for SPE.
Figure 8 (a) shows the spectra of 241Am calibration, representing NPEs from the COSINE-100 (black dashed line) and COSINE-100U (red solid line) encapsulation setups. The COSINE-100U setup provides approximately a 35% increase in light yield. From the mean of 59.54 keV peak, we observed a light yield of 19.6$\pm$0.3 NPE/keV for COSINE-100U C6, with similar improvements observed for all crystals, as summarized in Table 1.
Figure 8 (b) shows the 241Am calibration spectra in terms of energy. The COSINE-100U setup clearly demonstrates improved energy resolution compared to COSINE-100, benefiting from the increased light yield. Similar improvements were observed across all crystals.
@@
Figure 8: Measured spectra from the C6 crystal in the original COSINE-100 encapsulation and the newly designed COSINE-100U encapsulation at 59.54 keV using a 241Am source. (a) Comparison of light yield between COSINE-100 (black dashed line) and COSINE-100U (red solid line). (b) Comparison of energy resolution between COSINE-100 (black dashed line) and COSINE-100U (black dashed line).
@@
In the COSINE-100 experiment, C5 and C8 recorded relatively low light yields due to their 5-inch diameter optical windows, which were initially designed to accommodate 5-inch PMTs. This reduced light yield led to the exclusion of C5 and C8 from the low-energy dark matter search analysis. Additionally, C1 was excluded due to unexpected large noise events from its PMTs. Although the COSINE-100 experiment used eight crystals with a total mass of 106 kg, the effective detector mass used for the main physics analysis was only 61.4 kg Adhikari et al. (2021a, 2022a).
@@
Upon disassembling C5 and C8 for the COSINE-100U encapsulation, we observed liquid scintillator leakage inside the crystals through thin Mylar windows, which were initially used for low-energy x-ray calibration, such as the 5.9 keV signal from 55Fe. This leakage created grooves a few millimeters deep in the crystal surfaces. During the polishing process, we smoothed these grooves, but some remained, resulting in a reduced light yield of approximately 16.5 NPE/keV for C5 and C8. While this yield is lower than that of other COSINE-100U crystals, it is higher than the light yields of COSINE-100’s good-quality detectors, as summarized in Table 1. As a result, these two crystals can now be used for the physics analysis in the COSINE-100U experiment, increasing the effective detector mass to 99.1 kg.
Table 1: Summary of the COSINE-100 and COSINE-100U crystals, including their dimensions, mass, and light yields. The total mass of the COSINE-100 crystals was 106.3 kg, but only 61.4 kg of crystals were used for the dark matter search. Due to the machining process, the total mass of the COSINE-100U is reduced to 99.1 kg, with all crystals available for dark matter search analysis. Overall, the COSINE-100U setup provides a significantly increased light yield.
Crystal Size (inches) Mass [kg] Light yield [PEs/keV]
diameter×$\times$×length COSINE-100 COSINE-100U COSINE-100 COSINE-100U
C1 5.0$\times$7.0 8.3 7.1 14.9 $\pm$ 1.5 22.4 $\pm$ 0.5
C2 4.2$\times$11.0 9.2 8.7 14.6 $\pm$ 1.5 20.1 $\pm$ 0.5
C3 4.2$\times$11.0 9.2 8.7 15.5 $\pm$ 1.6 20.4 $\pm$ 0.4
C4 5.0$\times$15.3 18.0 16.9 14.9 $\pm$ 1.5 20.7 $\pm$ 0.4
C5 5.0$\times$15.5 18.3 17.2 7.3 $\pm$ 0.7 16.8 $\pm$ 0.5
C6 4.8$\times$11.8 12.5 11.7 14.6 $\pm$ 1.5 19.6 $\pm$ 0.3
C7 4.8$\times$11.8 12.5 11.6 14.0 $\pm$ 1.4 20.2 $\pm$ 0.5
C8 5.0$\times$15.5 18.3 17.2 3.5 $\pm$ 0.3 16.2 $\pm$ 0.4
Total 106.3 (61.4) 99.1

### v2

# Upgrading the COSINE-100 Experiment for Enhanced Sensitivity to Low-Mass Dark Matter Detection

@@ I Introduction
Numerous astronomical observations suggest that the majority of matter in the universe consists of invisible dark matter, though its nature and interactions remain elusive Clowe:2006eq ; Planck:2018vyg ; Bertone:2016nfn . Despite extensive efforts to directly detect dark matter, no definitive signals have been observed MarrodanUndagoitia:2015veg ; Schumann:2019eaa . The only exception is the DAMA/LIBRA experiment, which has reported annual modulation signals in a 250 kg array of NaI(Tl) detectors Bernabei:1998fta ; Bernabei:2013xsa ; Bernabei:2018yyw ; Bernabei:2021kdo , indicating potential dark matter-nuclei interactions Savage:2008er ; COSINE-100:2019brm . However, these results remain controversial, as no other experiments have observed similar signals Schumann:2019eaa ; ParticleDataGroup:2022pth , necessitating independent verification using the same NaI(Tl) crystal target materials.
@@
The COSINE-100 experiment concluded in March 2023 after 6.4 years of stable operation. It was relocated to the newly constructed Yemilab Park:2024sio ; Yemilab2024 in Korea for the next phase, COSINE-100U. During the transition, the crystal encapsulation was upgraded to improve light collection by directly attaching the crystals to photomultiplier tubes (PMTs) Choi:2020qcj . A similar technique, applied in the NEON experiment for measuring coherent elastic neutrino-nucleus scattering at nuclear reactors NEON:2022hbk , achieved approximately a 50% improvement in light yield and demonstrated long-term stability over two years Choi:2024trx . These improvements in light yield are expected to lower the energy threshold, thereby improving the sensitivity of COSINE-100U to dark matter, particularly in the low-mass region. For the COSINE-100U experiment at Yemilab, all eight NaI(Tl) crystals of COSINE-100 were upgraded. In this article, we detail the detector upgrade and the resulting improved sensitivities for low-mass dark matter detection.
@@ II NaI(Tl) Crystal Encapsulation
The COSINE-100 experiment Adhikari:2017esn was designed to investigate the DAMA/LIBRA annual modulation claim Bernabei:2021kdo using low-background NaI(Tl) detectors. The detector consists of eight NaI(Tl) crystals, each hermetically encapsulated in oxygen-free copper (OFC) enclosures and immersed in a liquid scintillator veto system Adhikari:2020asl to suppress external backgrounds. While COSINE-100 has successfully operated for over six years, improvements in light collection efficiency are necessary for the next phase, COSINE-100U, to enhance sensitivity to low-energy dark matter interactions.
One of the key upgrades in COSINE-100U involves a new crystal encapsulation technique to address the limitations of the original COSINE-100 design. The quartz windows used in the original setup introduced additional photon loss due to multiple optical interfaces between the optical pad, quartz window, and optical grease. The light collection efficiency was further limited by the mismatch between the crystal diameter and the PMTs, requiring beveled edges to guide photons toward the 3-inch PMTs. To improve light yield and signal quality, the COSINE-100U design eliminates the quartz windows and adopts a direct PMT-coupling approach, where the PMTs are attached directly to the crystal using a 2 mm thick optical pad. This method, initially demonstrated in the NEON experiment Choi:2020qcj ; Choi:2024trx , has been shown to increase light yield by approximately 50%, reaching approximately 22 photoelectrons (NPE)/keV, while maintaining long-term stability over two years.
NaI(Tl) crystals are typically packaged by commercial manufacurers using aluminum or copper enclosures with quartz windows, which protect the material while maintaining optical transparency. To enhance light collection efficiency, the crystal surfaces are often wrapped in reflective materials, such as Teflon sheets or aluminum oxide powder, which help maximize photon reflection. The COSINE-100 experiment utilized eight NaI(Tl) crystals grown by Alpha Spectra Inc., produced in collaboration with the KIMS, DM-Ice, and ANAIS collaborations to ensure low-background purity Kim:2014toa ; Amare:2014jta ; Adhikari:2015rba .
@@
Each of these cylindrical crystals, with slightly varying dimensions as listed in Table 1, was hermetically encased in OFC tubes (1.5 mm thick) with quartz windows at both ends, as exemplified by crystal-6 (C6) in Fig. 1. The lateral surfaces of each crystal were wrapped in Teflon sheets before being inserted into the OFC tubes. A 12 mm thick quartz window was coupled to the crystal using a 1.5 mm thick optical pad, and PMTs were attached to the quartz window using a small amount of high-viscosity optical grease.
@@
Figure 1: The encapsulation design of the COSINE-100 C6 crystal. A 4.8-inch diameter crystal is encased in OFC tubes, with an optical pad and quartz windows. The crystal shown in the figure is wrapped with a Teflon sheet. The 12 mm thick quartz window guides light from the 4.8-inch diameter crystal to the 3-inch PMTs, aided by a 45 angle surface. There are three optical interface layers between the crystal and the PMTs: a 1.5 mm optical pad, a 12 mm quartz windows, and a small amount of optical grease. The PMTs are encapsulated in OFC cases to protect them from the surrounding liquid scintillator.
@@
From the perspective of light collection efficiency, the original COSINE-100 encapsulation design had several drawbacks. The three-layer optical interface, comprising the optical pad, quartz window, and optical grease, introduced additional reflections, reducing light collection efficiency. Although the 12 mm thick quartz window, with a 45 angle, guided photons from the 4.8-inch diameter crystal, it was insufficient for efficiently directing photons to the 3-inch PMTs. As shown in Fig. 1, uncovered areas resulted in photon reflection, further reducing light collection efficiency.
@@
To address these issues, an enhanced crystal encapsulation technique was developed, initially in the NEON experiment, which involved directly attaching the PMT to the crystal using only a 2 mm thick optical pad Choi:2020qcj . In this method, the crystal’s diameter was matched to the 3-inch diameter of the PMTs, eliminating photon absorption in the quartz and minimized photon loss due to reflections from multiple interfaces. This modification resulted in a light yield of up to 22 NPE/keV, approximately 50% higher than the light yield observed in COSINE-100 crystals, which are measured around 15 NPE/keV Adhikari:2017esn .
@@
This technique was initially applied in the NEON experiment during an engineering run at a nuclear reactor to observe coherent elastic neutrino-nucleus scattering NEON:2022hbk . However, some design weaknesses were identified, as liquid scintillator leakage into the detector caused a gradual decrease in crystal light yield and increase of PMT-induced noise.
@@
The updated design separates the encapsulation into two components: an inner structure to maintain a stable coupling between the PMTs and crystal, and an outer OFC case to prevent the infiltration of external air and liquid scintillator. These improvements were successfully applied to the NEON experiment, which has been collecting stable physics data for over two years with a light yield exceeding 22 NPE/keV Choi:2024trx .
Figure 2: The encapsulation design of the COSINE-100U C6 crystal. The edges of a 4.8-inch diameter crystal are beveled at a 45 angle to guide light to 3-inch PMTs and wrapped with a Teflon sheet. PMTs are directly attached to the crystal using a 2 mm thick optical pad and supported by a rigid 5 mm thick PTFE inner structure. The crystal-PMT assembly, along with the inner structure, is encased in OFC tube to protect it from the surrounding liquid scintillator.
Based on this experience, some of the current authors, who were also involved in the NEON experiment, applied the lessons learned to improve the crystal encapsulation for COSINE-100U, as shown in Fig. 2. While the NEON crystals have a 3-inch diameter matching the size of 3-inch PMTs, the COSINE-100 crystals have larger diameters, as listed in Table 1. To address this, the edges of the COSINE-100 crystals are beveled at a 45 angle, reducing the edge diameter to 3-inch and effectively guiding light from the larger-diameter crystals to the 3-inch PMTs.
To ensure stable mounting of the PMTs to the crystals, a 5 mm thick PTFE inner structure is employed, which fully encases the crystal. A 2 mm thick optical pad is positioned between the PMTs and the crystal surface to optimize light transmission. Stable light coupling is achieved through a PTFE ring, secured with brass bolts to the PTFE inner structure, which applies consistent and uniform pressure to the optical pad. The assembled crystals are hermetically encased in 2 mm thick OFC tubes, with both ends sealed by 20 mm thick OFC lid flanges. Further details regarding the COSINE-100U crystal assembly process can be found in Appendix I.
Table 1: Summary of the COSINE-100 and COSINE-100U crystals geometry. The COSINE-100 experiment originally consisted of eight NaI(Tl) crystals with a total mass of 106.3 kg. However, three crystals (marked with an asterisk) were excluded from the dark matter search, resulting in an effective mass of 61.4 kg. In the COSINE-100U, all eight crystals were included in the analysis, but the total mass was reduced to 99.1 kg due to the machining process.
Crystal Size (inches) Mass [kg]
diameter×$\times$×length COSINE-100 COSINE-100U
C1 5.0$\times$7.0 8.3 7.1
C2 4.2$\times$11.0 9.2 8.7
C3 4.2$\times$11.0 9.2 8.7
C4 5.0$\times$15.3 18.0 16.9
C5 5.0$\times$15.5 18.3 17.2
C6 4.8$\times$11.8 12.5 11.7
C7 4.8$\times$11.8 12.5 11.6
C8 5.0$\times$15.5 18.3 17.2
Total 106.3 (61.4) 99.1
@@ III.1 Measurement Setup
Upon assembling the COSINE-100 crystals, the Yemilab facility was not yet ready for operation of the COSINE-100U experiment. For the initial crystal characterization, we employed a simple shielding setup at sea level in the experimental hall of the Institute for Basic Science (IBS) in Korea. This setup consisted of two layers of shielding: 10 cm thick lead and 20 cm thick liquid scintillator, which also functioned as an active veto detector. The schematic view of the shielding structure is shown in Fig. 3(a). The liquid scintillator was housed within a 124.5 cm $\times$ 49.5 cm $\times$ 49.5 cm cubic stainless steel box. Three 8-inch PMTs were used to read the signals from the liquid scintillator as shown in Fig. 3(b). This setup was initially developed as a prototype detector for the NEOS experiment NEOS:2016wee and was reused for this test. Inside the container, an acrylic table was used to install one NaI(Tl) crystal for the initial evaluation of its performance, as shown in Fig. 3(b).
Figure 3: Test measurements setup. (a) Schematic view of the shield structure used for the sea-level measurement at the Institute for Basic Science. The encapsulated crystal detector is installed within the inner liquid scintillator layer, approximately 20 cm thick. This layer is surrounded by a 10 cm thick lead shield and an outer liquid scintillator container. For this measurement, the outer liquid scintillator container was not filled. (b) The C1 crystal is installed in this setup for light yield and stability measurements.
Events were collected by two PMTs attached to each crystal, with the photoelectron signals amplified using a preamplifier and digitized by a flash analog-to-digital converter (FADC) at a sampling rate of 500 MHz. The FADC, connected to the trigger control board (TCB), recorded events that satisfied specific trigger conditions, capturing waveforms over an 8 $\mu$s window starting 2.4 $\mu$s before the trigger time COSINE-100:2018rxe . Additionally, signals from three 8-inch PMTs for the liquid scintillator were recorded using another 500 MHz FADC with the same 8 $\mu$s recording length when NaI(Tl) crystal-triggered events occurred.
@@ III.2 Light Yield Measurement
We measured the light yield of the newly encapsulated COSINE-100U crystal by irradiating it with 59.54 keV $\gamma$-rays from a 241Am source. The 241Am source was positioned above the crystal in Fig. 3(b). The mean charge corresponding to a single photoelectron (SPE) was determined by analyzing trailing isolated cluster pulses in two specific time windows (5–7 and 6–8 $\mu$s) within 8 $\mu$s long waveforms. These windows, located 2.6–4.6 and 3.6–5.6 $\mu$s from the triggering position, were chosen to minimize the impact from large photoelectron clusters. We simultaneously fit the 5–7 and 6–8 $\mu$s clusters using models that included up to four photoelectrons clusters Choi:2024ziz ; Choi:2024trx , as shown in Fig. 4. The number of photoelectrons (NPE) was calculated by dividing the integrated charge of the main pulse (within 5 $\mu$s of the pulse start) by the charge of SPE.
@@
Figure 4: Estimation of mean charge of single photoelectrons (SPEs). The charge distribution of the isolated clusters in the 5–7 $\mu$s (red dots) and 6–8 $\mu$s (blue dots) windows are modeled with up to four NPE clusters, simultaneously fitted for both the 5–7 and 6–8 $\mu$s windows, yielding a mean charge of 597.2 ADC for SPE.
[…]
Figure 5(b) shows the 241Am calibration spectra in terms of energy, with fit results using a Crystal Ball function to account for the asymmetric shape of 59.54 keV peak. The asymmetric shape of the lower-energy shoulder is attributed to partial energy deposition in the surrounding encapsulation material due to Compton scattering. For the COSINE-100U C6 crystal, the improved light yield resulted in an approximately 6% better energy resolution compared to the COSINE-100 setup. Similar improvements were observed for all crystals, as summarized in Table 2.
@@
Figure 5: Light yield improvement of the COSINE-100U encapsulation. Measured spectra from the C6 crystal in the original COSINE-100 encapsulation and the newly designed COSINE-100U encapsulation at 59.54 keV using a 241Am source are compared. (a) Comparison of light yield between COSINE-100 (black dashed line) and COSINE-100U (red solid line). (b) Comparison of energy resolution at 59.54 keV between the COSINE-100 (black solid line) and COSINE-100U (red solid line) setups. The data have been modeled using Crystal Ball functions to account low energy shoulder caused by partial energy deposition in the encapsulation material.
@@
In the COSINE-100 experiment, C5 and C8 recorded relatively low light yields due to their 5-inch diameter optical windows, which were initially designed to accommodate 5-inch PMTs. This reduced light yield led to the exclusion of C5 and C8 from the low-energy dark matter search analysis. Additionally, C1 was excluded due to unexpected large noise events from its PMTs. Although the COSINE-100 experiment included eight crystals with a total mass of 106 kg, the effective detector mass used for the main physics analysis was only 61.4 kg COSINE-100:2021xqn ; COSINE-100:2021zqh .
@@
Upon disassembling C5 and C8 for the COSINE-100U encapsulation, we observed liquid scintillator leakage inside the crystals through thin Mylar windows, which were initially used for low-energy x-ray calibration, such as the 5.9 keV signal from 55Fe. This leakage created grooves a few millimeters deep in the crystal surfaces. During the polishing process, we smoothed these grooves, but some remained, resulting in a reduced light yield of approximately 16.5 NPE/keV for C5 and C8. While this yield is lower than that of other COSINE-100U crystals, it is higher than the light yields of COSINE-100’s good-quality detectors, as summarized in Table 2. As a result, these two crystals can now be used for the physics analysis in the COSINE-100U experiment. In addition, C1 was recovered by replacement its PMT, which successfully eliminated the noise issue, thereby increasing the effective detector mass to 99.1 kg.
Table 2: Summary of the COSINE-100U energy resolution and light yields. The root-mean-square (RMS) energy resolution at the 59.54 keV $\gamma$ peak, determined using a 241Am source calibration, and the measured light yields of the crystals are compared between the COSINE-100 and COSINE-100U setups. The COSINE-100U setup demonstrates a significantly improved energy resolution, attributed to the increased light yield achieved through the upgraded design and enhancements in light collection efficiency.
Crystal Light yield [NPE/keV] 59.54 keV RMS resolution [%]
COSINE-100 COSINE-100U COSINE-100 COSINE-100U
C1 14.9 $\pm$ 1.5 22.4 $\pm$ 0.5 4.80$\pm$0.06 4.41$\pm$0.04
C2 14.6 $\pm$ 1.5 20.1 $\pm$ 0.5 5.06$\pm$0.06 4.54$\pm$0.04
C3 15.5 $\pm$ 1.6 20.4 $\pm$ 0.4 4.96$\pm$0.05 4.42$\pm$0.04
C4 14.9 $\pm$ 1.5 20.7 $\pm$ 0.4 4.80$\pm$0.05 4.59$\pm$0.03
C5 7.3 $\pm$ 0.7 16.8 $\pm$ 0.5 8.33$\pm$0.11 5.04$\pm$0.03
C6 14.6 $\pm$ 1.5 19.6 $\pm$ 0.3 4.56$\pm$0.05 4.33$\pm$0.02
C7 14.0 $\pm$ 1.4 20.2 $\pm$ 0.5 4.86$\pm$0.04 4.53$\pm$0.03
C8 3.5 $\pm$ 0.3 16.2 $\pm$ 0.4 11.53$\pm$0.11 5.06$\pm$0.04